# Visual Haystacks: A Vision-Centric Needle-In-A-Haystack Benchmark

### v2

@@ 2 The Visual Haystacks Benchmark (VHs)
We address these shortcomings in the Visual Haystacks (VHs) benchmark which emphasizes realistic, visually grounded needles and mirrors real-world visual long-context learning tasks. The design of VHs is based on two core principles: realism and reliability. The needles in VHs are real objects within natural images, overcoming the artificial inclusion of overlaid text or image patches (Reid et al., 2024; Wang et al., 2024b) that previous visual NIAH benchmarks suffer from. VHs utilizes in-domain images and straightforward questions with human-annotated ground truth. Compared to recent LLM and LMM datasets (Gema et al., 2024; Zhu et al., 2024) which face noises, biases, and out-of-distribution (OOD) challenges due to their data labeling methodology, VHs is much more reliable. Finally, VH is diverse and large-scale; haystack sizes are as large as 10,000 images, haystacks can feature one or more needles, and each needle can draw from over 50 distinct objects. With over 1,000 examples of each setting, VHs provides 97k images total for comprehensive visual NIAH evaluation.
@@ 2.1 Benchmark Construction
After selecting the object pairs, setting the questions, and determining the answers, we then compile corresponding image haystacks by first curating the needle images, which contain the “anchor object” as indicated by the paired answer. We then accompany these images with multiple negative distractors to form haystacks of varying sizes, ranging from 1 to 10,000 haystack images. We select distractor images in a manner such that no distrator contains any anchor object (following COCO’s object annotations). However, some of them contain the target objects so as to create meaningful distractors. The single-needle setting contains only one needle image in the haystack, while the multi-needle setting includes either two and three needle images. VHs consists of 1000 question-answer pairs for both single- and multi-needle settings, with an explicit small subset $\text{VHs}_{\text{small}}$ consisting of 100 questions, which is helpful for economical evaluation of expensive closed-source models. Appendix A illustrates several examples and statistics of single- and multi-needle benchmarks.
@@ 3 Can LMMs handle Long-Context Visual Inputs?
We evaluated a range of state-of-the-art open-source and proprietary LMMs on the VHs benchmark. Specifically, this includes GPT-4o (OpenAI, 2024), Gemini 1.5 Pro (Reid et al., 2024), Qwen2-VL (Wang et al., 2024a), Phi-3 Vision (Abdin et al., 2024), mPLUG-Owl3 (Ye et al., 2024), and InternVL2 (Chen et al., 2024).
We also implement two non-LMM baselines to further contextualize LMM performance. The first baseline, termed “caption aggregation,” follows a two-stage process in which images are first captioned using Qwen2-VL, and subsequently, Llama-v3.1 (Dubey et al., 2024) is used to answer questions based on the aggregated captions. This is a query-unaware baseline given that this method generates captions without conditioning on the questions being asked, capturing a sub-optimal, yet informative, comparison point. The second baseline, the “detector oracle,” serves as an upper-bound baseline by utilizing OWLv2 (Minderer et al., 2024) as an object detection model to first find all images with the anchor object which are then each checked for the target object. Detailed descriptions of both baselines are available in subsection B.1.
In both single-needle and multi-needle settings, we conducted experiments using the full VHs dataset where the haystack size was 100 images or fewer, and switched to the $\text{VHs}_{\text{small}}$ subset with larger haystacks to mitigate computational costs. We report the results as bootstrapped averages, with standard deviations indicated as error bars to ensure robustness. The outcomes of the single- and multi-needle challenges are presented in Figure 2 and Figure 3. Figure 4 explores LMM behavior when the needle image is located in different positions within the input image set. Comprehensive results are shown in subsection B.2.
@@ 3.2 Difficulty in Reasoning Across Multiple Images
In the full VHs multi-needle challenge, where models must both retrieve visual information and reason across multiple images, the results (shown in Figure 3 (B)) were consistent with those from the single-needle experiments. As the number of input images increased, performance steadily declined across all models. While Gemini v1.5 Pro maintained relatively high accuracy (above 65%, with only a 5% decline from N=5 to N=100), other models experienced steeper drops, with some falling below the caption aggregation baseline. It is particularly noteworthy that GPT-4o, which performed best in the oracle multi-needle experiments with two or three needles, struggled significantly in real-world multi-needle challenges with distractor images, performing worse than Gemini v1.5 Pro by a large margin. Open-source models also demonstrated performance declines from the oracle scenario to the real-world multi-needle challenges, with mPLUG-Owl3 dropping from the top-performing open-source model to the worst among its peers. Overall, these results highlight the persistent challenges that LMMs face in managing complex, multi-image visual retrieval and reasoning tasks.
@@ 3.3 Positional Biases
As shown in Figure 4, LMMs are also sensitive to the positional placement of the needle image. Interestingly, this positional bias varies across models. For example, Gemini v1.5 Pro tends to favor images at the beginning of the sequence, while GPT-4o exhibits a clear lost-in-the-middle effect, similar to what has been observed in natural language tasks (Liu et al., 2024b). In contrast, open-source models like Qwen2-VL, Phi-3-vision, and InternVL2 favor images closest to the question (i.e., the last image in the sequence) given a small number of images ($N\leq 5$). We break down the 2D heatmap for each method in Appendix B.
@@ 4 MIRAGE: Multi-Image Retrieval Augmented Generation
Despite its simplicity, MIRAGE is the first framework to enable visual-RAG operations with LMMs. MIRAGE can take over 10,000 images during inference with a single 40GB A100 GPU while maintaining competitive performance across various tasks, as demonstrated in section 5.
@@ 5 Results & Discussion
Method RetVQA (MIQA) VQAv2 GQA TextVQA POPE MMB MMB-CN MM-Vet
GPT-4o (OpenAI, 2024) 34.6 77.2 - 78.0 - - - -
Gemini-v1.5 (Reid et al., 2024) 32.2 73.2 - 73.5 - - - -
LWM (Liu et al., 2024a) - 55.8 44.8 18.8 75.2 - - 9.6
LLaVA-v1.5-7B Liu et al. (2023a) 30.6 78.5 62.0 58.2 85.9 64.3 58.3 31.1
MIRAGE-8.3B (Ours) 67.6 76.6 59.1 56.2 85.4 69.2 66.9 33.4
Table 1: Comparative performance of methods on multi-image and single-image QA tasks. MIRAGE shows strong performance in multi-image QA and competitive performance in single-image QA compared to existing proprietary and open-source long-context models. Built primarily on LLaVA-v1.5-7B, MIRAGE-8.3B matches its performance on single-image QA while excelling in multi-image QA, achieving this with a reduced number of tokens per image.
Method Tokens/Img VQAv2 GQA Vizwiz TextVQA POPE MMB MMB-CN MM-Vet
Original LLaVA 576 78.5 62.0 50.0 58.2 85.9 64.3 58.3 31.1
3x3 Max-Pooling 64 68.7 56.2 41.3 48.5 83.0 59.2 49.3 24.3
Global Avg. Pooling 1 62.5 51.3 37.7 45.5 79.6 55.0 45.5 18.9
Q-Former (Ours) 32 72.8 56.6 48.0 47.1 83.9 61.5 55.0 27.3
Table 2: Exploration of various token reduction methods. The Q-former demonstrates the most efficiency in reducing token count while maintaining the majority of general QA performance. All experiments were conducted using the official LLaVA-v1.5-7B model with Vicuna-v1.5 as the LLM and LLaVA’s instruction tuning data. Note that the Q-former variant shows lower performance compared to our final model MIRAGE-8.3B, which leverages better training data and a more powerful LLM, Llama-v3.1-8B.
Visual Haystacks: MIRAGE’s performance on single- and multi-needle VHs challenge is presented in Figure 2 and Figure 3. All baselines and MIRAGE are tested in an identical, zero-shot setting in which they are not trained on these kinds of questions. In the single-needle challenge, MIRAGE demonstrates competitive oracle (N=1) performance against existing approaches. Notably, MIRAGE is the only method that scales to 10k input images, outperforming all open-source models when processing more than three images, and surpassing both Gemini v1.5 Pro and GPT-4o when using over 50 images.
@@
Conventional VQA Tasks: We also compare MIRAGE with other methods on conventional VQA tasks, including the RetVQA multi-image task (Penamakuri et al., 2023) and several single-image QA tasks, VQA-v2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), TextVQA (Singh et al., 2019), POPE (Li et al., 2023b), MMB (Liu et al., 2024c), MMB-CN (Liu et al., 2024c), and MM-Vet (Yu et al., 2023b). In all cases, we follow standard evaluation procedures from either RetVQA (Penamakuri et al., 2023) or LLaVA (Liu et al., 2023b). As shown in Table 1, MIRAGE is the only method capable of excelling in both multi-image and single-image QA tasks. Open-source models like LLaVA-v1.5-7B (Liu et al., 2023a) struggle with multi-image QA, likely due to insufficient context length and lack of specific training for this task. Similarly, commercial long-context LMMs also perform poorly on multi-image tasks containing a lot of distracted images.
Image Token Compression Scheme: MIRAGE aims to facilitate future research into MIQA tasks, with a critical focus on reducing the context length required per image. To investigate this, we compared various token compression schemes in the LLaVA-v1.5-7B model. These schemes included max-pooling and average-pooling compression, reducing an image to a single token. All models were trained using LLaVA’s single-image instruction fine-tuning dataset, with the retriever disabled for this experiment. The results, presented in Table 2, indicate that while previous work has suggested that Q-Former is sub-optimal (Fan et al., 2024), our findings show that Q-Former retains reasonable performance while significantly reducing context length.
@@
Figure 6: Comparison between MIRAGE retriever and CLIP retriever across both performance and runtime. We can see that while CLIP is slightly faster, MIRAGE has significantly higher recall, which impacts downstream performance. Both MIRAGE approaches with a retrieval component are significantly more efficient than GPT-4o and Caption-based baselines.
@@ 6 Related Work
Ours is not the first work to address the MIQA problem. Similar to our work, Bansal et al. (Bansal et al., 2020) introduce a method for multi-image question answering, however in their approach, all of the images (up to 10) are relevant to the question, and their method does not contain a retrieval component. Penamakur et al. (Penamakuri et al., 2023) propose a dataset and method for retrieval-based visual question answering (RetVQA), and share with our work the fact that answers must be gleaned from a set of both relevant and irrelevant images. However, their work differs from ours in two key ways: (1) RetVQA question contexts contain at most two relevant images, meaning that models do not have to reason across many images, largely sidestepping the problem of limited context length, and (2) their approach only allows for small image worlds (up to 30), meaning that they can use a pairwise encoder for each image, and they do not need to search over a large dataset of images potentially containing distractors, limiting the efficiency of their approach. Finally, they are only able to perform question answering over the images, whereas we pursue a method that can additionally perform more complex reasoning tasks given the images in the dataset.
@@
Similar to our MIRAGE approach is Chen et al. (Chen et al., 2022) and Yasunaga et al. (Yasunaga et al., 2022) which both retrieve multiple visual documents to answer queries. In the case of Chen et al. (Chen et al., 2022), the queries are open-ended question answering, and thus, are not grounded within a particular set of context images. Yasunaga et al. (Yasunaga et al., 2022) focuses on one and few-shot classification and image generation, and does not use their multi-image retrieval to answer aggregated questions about those images. In single-image QA, image retrieval in large multimodal models (LMMs) has been explored using “retrieval-tokens” (Koh et al., 2023), however, it is unclear how such an approach would scale to multi-image QA problems with multiple relevant images. Complementary to our approach, several methods have focused on true long-context LMMs by introducing down-sampling techniques or state-space models (Li et al., 2023a; Chen et al., 2024; Abdin et al., 2024; Wang et al., 2024c) - and while these methods cannot yet handle 10K images well, we expect in the future for this to be a promising direction of further research. In addition to general QA, several other works contain domain-specific multi-image QA problems, including Slide VQA (Tanaka et al., 2023), Multimodal QA (Talmor et al., 2021), WebQA (Chang et al., 2022) and Document VQA (Tito et al., 2021), however, none are explicitly designed to answer integration questions across multiple images from a subset of large, unrelated images (like VHs).
@@
Outside of VQA, In traditional NLP, retrieving small passages or single documents from large-scale corpuses has proven effective. Zhang et al. (Zhang et al., 2024) introduces a method that fine-tunes LLMs on both relevant and irrelevant documents to support better RAG performance, but does not train an explicit query-aware filter or compression module. ATLAS (Izacard et al., 2023) treats documents as latent variables during training, allowing for efficient retrieval, but requires a complex joint-training setup. Similarly, several methods (Shi et al., 2023; Ram et al., 2023; Borgeaud et al., 2022; Khandelwal et al., 2019) have demonstrated success with zero or few-shot augmentation of standard LLM contexts with retrieved documents. Beyond context augmentation, several traditional NLP approaches (Lin et al., 2023b; Wang et al., 2023; Xu et al., 2023; Liu et al., 2024d; Xiang et al., 2024; Xu et al., 2024) have demonstrated that fine-tuning LLMs to be robust to noisy RAG outputs can lead to performance improvements.
Contemporary to VHs, Wang et al. (2024b) introduce a synthetic benchmark to evaluate LMMs on multi-image question answering (up to 36 images per haystack). Similarly, Zhao et al. (2024) propose a video-centric benchmark to test video LMMs on spatiotemporal understanding. Unlike VHs which relies on object labels in real images to build questions about up to 10K images, both benchmarks focus heavily on synthetic content with short-range retrieval problems (<100 frames/images). Moreover, similar to traditional visual NIAH benchmarks, they incorporate out-of-distribution content (frames, patches) into the image or video as the visual “needle.” This contrasts sharply with VHs, which employs realistic and in-distribution visual content, thereby posing a more complex challenge that is less susceptible to straightforward solutions.

### v3

@@ 2 The Visual Haystacks Benchmark (VHs)
We address these shortcomings in the Visual Haystacks (VHs) benchmark which emphasizes realistic, visually grounded needles and mirrors real-world visual long-context learning tasks. The design of VHs is based on two core principles: realism and reliability. The needles in VHs are real objects within natural images, overcoming the artificial inclusion of overlaid text or image patches (Reid et al., 2024; Wang et al., 2024c) that previous visual NIAH benchmarks suffer from. VHs utilizes in-domain images and straightforward questions with human-annotated ground truth. Compared to recent LLM and LMM datasets (Gema et al., 2024; Zhu et al., 2024) which face noises, biases, and out-of-distribution (OOD) challenges due to their data labeling methodology, VHs is much more reliable. Finally, VH is diverse and large-scale; haystack sizes are as large as 10,000 images, haystacks can feature one or more needles, and each needle can draw from over 50 distinct objects. With over 1,000 examples of each setting, VHs provides 97k images total for comprehensive visual NIAH evaluation.
@@ 2.1 Benchmark Construction
After selecting the object pairs, setting the questions, and determining the answers, we then compile corresponding image haystacks by first curating the needle images, which contain the “anchor object” as indicated by the paired answer. We then accompany these images with multiple negative distractors to form haystacks of varying sizes, ranging from 1 to 10,000 haystack images. We select distractor images in a manner such that no distrator contains any anchor object (following COCO’s object annotations). However, some of them contain the target objects so as to create meaningful distractors. The single-needle setting contains only one needle image in the haystack, while the multi-needle setting includes either two and three needle images. VHs consists of 1000 question-answer pairs for both single- and multi-needle settings, with an explicit small subset $\text{VHs}_{\text{small}}$ consisting of 100 questions, which is helpful for economical evaluation of expensive closed-source models. Appendix B illustrates several examples and statistics of single- and multi-needle benchmarks.
@@ 3 Can LMMs handle Long-Context Visual Inputs?
We evaluated a range of state-of-the-art open-source and proprietary LMMs on the VHs benchmark. Specifically, this includes GPT-4o (OpenAI, 2024), Gemini 1.5 Pro (Reid et al., 2024), Qwen2-VL (Wang et al., 2024b), Phi-3 Vision (Abdin et al., 2024), mPLUG-Owl3 (Ye et al., 2024), and InternVL2 (Chen et al., 2024).
We implement two non-LMM baseline approaches to further contextualize LMM performance. The first baseline, termed “caption aggregation,” follows a two-stage process in which images are first captioned using Qwen2-VL, and subsequently, Llama-v3.1 (Dubey et al., 2024) is used to answer questions based on the aggregated captions. This is a query-unaware baseline given that this method generates captions without conditioning on the questions being asked, capturing a sub-optimal, yet informative, comparison point. The second baseline, the “detector oracle,” serves as an upper-bound baseline by utilizing OWLv2 (Minderer et al., 2024) as an object detection model to first find all images with the anchor object which are then each checked for the target object. Detailed descriptions of both baselines are available in subsection C.1.
In both single-needle and multi-needle settings, we conducted experiments using the full VHs dataset where the haystack size was 100 images or fewer, and switched to the $\text{VHs}_{\text{small}}$ subset with larger haystacks to mitigate computational costs. We report the results as bootstrapped averages, with standard deviations indicated as error bars to ensure robustness. The outcomes of the single- and multi-needle challenges are presented in Figure 2 and Figure 3. Figure 4 explores LMM behavior when the needle image is located in different positions within the input image set. Comprehensive results are shown in subsection C.2.
@@ 3.2 Difficulty in Reasoning Across Multiple Images
In the full VHs multi-needle challenge, where models must both retrieve visual information and reason across multiple images, the results (shown in Figure 3 (B)) were consistent with those from the single-needle experiments. As the number of input images increased, performance steadily declined across all models. While Gemini v1.5 Pro maintained relatively high accuracy (above 65%, with only a 5% decline from N=5 to N=100), other models experienced steeper drops, with some falling below the caption aggregation baseline. It is particularly noteworthy that GPT-4o, which performed best in the oracle multi-needle experiments with two or three needles, struggled significantly in real-world multi-needle challenges with distractor images, performing worse than Gemini v1.5 Pro by a large margin. Open-source models also demonstrated performance declines from the oracle scenario to the real-world multi-needle challenges, with mPLUG-Owl3 dropping from the top-performing open-source model to the worst among its peers. Overall, these results highlight the persistent challenges that LMMs face in managing complex, multi-image visual retrieval and reasoning tasks. More in-depth analyses are reported in subsection C.4.
@@ 3.3 Positional Biases
As shown in Figure 4, LMMs are also sensitive to the positional placement of the needle image. Interestingly, this positional bias varies across models. For example, Gemini v1.5 Pro tends to favor images at the beginning of the sequence, while GPT-4o exhibits a clear lost-in-the-middle effect, similar to what has been observed in natural language tasks (Liu et al., 2024b). In contrast, open-source models like Qwen2-VL, Phi-3-vision, and InternVL2 favor images closest to the question (i.e., the last image in the sequence) given a small number of images ($N\leq 5$). We break down the 2D heatmap for each method in Appendix C.
@@ 4 MIRAGE: Multi-Image Retrieval Augmented Generation
Despite its simplicity, MIRAGE is the first general framework to enable visual-RAG operations with LMMs. The collaboration between a retriever and a processor plays a crucial role in real-world, large-scale multi-image QA scenarios, such as photo album searches or geospatial image analysis. By filtering images based on relevance before reasoning occurs in the LMM, the retriever ensures efficient runtime (see Figure 6) and prevents the LMM from being affected by visual distractors—a known limitation of current LMMs. Furthermore, the LMM, trained with our custom dataset, ensures competitive performance across diverse and general question-answering tasks, as demonstrated in section 5.
@@ 5 Results & Discussion
Method RetVQA VQAv2 GQA TextVQA POPE MMB MMB-CN MME SEED MM-Vet
GPT-4o 34.6 77.2 - 78.0 87.2 - - 1614.2 - -
Gemini v1.5 Pro 32.2 73.2 - 73.5 88.2 - - 1562.4 - -
LLaVA-v1.5-7B 30.6 78.5 62.0 58.2 85.9 64.3 58.3 1510.7 58.6 31.1
LWM - 55.8 44.8 18.8 75.2 - - - - 9.6
MIRAGE-8.3B (Ours) 67.6 76.6 59.1 56.2 85.4 69.2 66.9 1437.9 59.0 33.4
Table 1: Comparative performance of methods on multi-image and single-image QA tasks. Performance metrics for single-image QA baselines are sourced from their official papers or the online leaderboard. MIRAGE shows strong performance in multi-image QA and competitive performance in single-image QA compared to existing proprietary and open-source long-context models. Built primarily on LLaVA-v1.5-7B, MIRAGE-8.3B matches its performance on single-image QA while excelling in multi-image QA, achieving this with an 18x reduction in tokens per image.
Visual Haystacks: MIRAGE’s performance on single- and multi-needle VHs challenge is presented in Figure 2 and Figure 3. All baselines and MIRAGE are tested in an identical, zero-shot setting in which they are not trained on these kinds of questions. While MIRAGE operates as a visual-RAG framework with a retriever and an LMM, it is fundamentally an end-to-end trained “single-model architecture” (Figure 5). In this context, the “retriever” component should be understood as a component for filtering irrelevant images and enhancing long-context reasoning. Based on this, we believe that our comparison with other “single-model LMMs” on VHs is both fair and valid. In the single-needle challenge, MIRAGE demonstrates competitive oracle (N=1) performance against existing approaches. Notably, MIRAGE is the only method that scales to 10k input images, outperforming all open-source models when processing more than three images, and surpassing both Gemini v1.5 Pro and GPT-4o when using over 50 images.
@@
Conventional VQA Tasks: Although the primary focus of this paper is on large-scale multi-image QA tasks like VHs, we also evaluate MIRAGE on conventional VQA tasks, including the RetVQA multi-image QA (Penamakuri et al., 2023) and several single-image QA tasks, such as VQA-v2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), TextVQA (Singh et al., 2019), POPE (Li et al., 2023c), MMB (Liu et al., 2024c), MMB-CN (Liu et al., 2024c), MME (Fu et al., 2023), SEED-Bench (Li et al., 2023a), and MM-Vet (Yu et al., 2023b). Baseline comparisons include GPT-4o (OpenAI et al., 2024), Gemini v1.5 Pro (Reid et al., 2024), LWM (Liu et al., 2024a), and LLaVA-v1.5-7B (Liu et al., 2023a). In all cases, we follow standard evaluation procedures from either RetVQA (Penamakuri et al., 2023) or LLaVA (Liu et al., 2023b). As shown in Table 1, MIRAGE achieves state-of-the-art performance on the RetVQA multi-image QA benchmark. On single-image QA tasks, MIRAGE exhibits minor performance degradation on VQA-v2, GQA, POPE, and TextVQA compared to LLaVA-v1.5-7B, which we attribute to MIRAGE’s token compression scheme (See subsection C.3 for more discussions). However, MIRAGE surpasses LLaVA-v1.5-7B on benchmarks like MMB, MMB-CN, and MM-Vet, consistently outperforming LWM across all evaluated tasks. Additionally, MIRAGE remains competitive with proprietary models, making it one of the most versatile open-source frameworks for both single- and multi-image QA tasks.
@@
Conclusion: The experiments highlight three primary advantages of MIRAGE: Scalability, as it is the only framework capable of processing thousands of images in a single query, significantly exceeding the capacity of existing models; Performance, with state-of-the-art results on multi-image tasks and competitive results on single-image tasks despite the usage of fewer tokens per image; and Efficiency, achieved through token reduction via the retriever component and the co-trained Q-former compressor, allowing MIRAGE to run on a single 40GB A100 GPU with faster runtimes and better retrieval performance compared to CLIP.
Figure 6: Analyses on MIRAGE and CLIP retriever across performance and runtime. Both approaches with a retrieval component are significantly more efficient than GPT-4o and caption-based baselines. While CLIP is slightly faster, MIRAGE has significantly higher recall, which impacts downstream performance.
@@ 6 Related Work
Ours is not the first work to address the MIQA problem. Similar to our work, Bansal et al. (2020) introduce a method for multi-image question answering, however in their approach, all of the images (up to 10) are relevant to the question, and their method does not contain a retrieval component. Penamakuri et al. (2023) propose a dataset and method for retrieval-based visual question answering (RetVQA), and share with our work the fact that answers must be gleaned from a set of both relevant and irrelevant images. However, their work differs from ours in two key ways: (1) RetVQA question contexts contain at most two relevant images, meaning that models do not have to reason across many images, largely sidestepping the problem of limited context length, and (2) their approach only allows for small image worlds (up to 30), meaning that they can use a pairwise encoder for each image, and they do not need to search over a large dataset of images potentially containing distractors, limiting the efficiency of their approach. Finally, they are only able to perform question-answering over the images, whereas we pursue a method that can additionally perform more complex reasoning tasks given the images in the dataset.
@@
Similar to our MIRAGE approach are Chen et al. (2022) and Yasunaga et al. (2022) which both retrieve multiple visual documents to answer queries. In the case of Chen et al. (2022), the queries are open-ended question answering, and thus, are not grounded within a particular set of context images. Yasunaga et al. (2022) focuses on one and few-shot classification and image generation, and does not use their multi-image retrieval to answer aggregated questions about those images. In single-image QA, image retrieval in large multimodal models (LMMs) has been explored using “retrieval-tokens” (Koh et al., 2023), however, it is unclear how such an approach would scale to multi-image QA problems with multiple relevant images. Complementary to our approach, several methods have focused on true long-context LMMs by introducing down-sampling techniques or state-space models (Gupta et al., 2024; Li et al., 2023b; Chen et al., 2024; Abdin et al., 2024; Wang et al., 2024d) - and while these methods cannot yet handle 10K images well, we expect in the future for this to be a promising direction of further research. In addition to general QA, several other works contain domain-specific multi-image QA problems, including Slide VQA (Tanaka et al., 2023), Multimodal QA (Talmor et al., 2021), WebQA (Chang et al., 2022) and Document VQA (Tito et al., 2021).
@@
Outside of VQA, In traditional NLP, retrieving small passages or single documents from large-scale corpora has proven effective. Zhang et al. (2024) introduces a method that fine-tunes LLMs on both relevant and irrelevant documents to support better RAG performance, but does not train an explicit query-aware filter or compression module. ATLAS (Izacard et al., 2023) treats documents as latent variables during training, allowing for efficient retrieval, but requires a complex joint-training setup. Similarly, several methods (Shi et al., 2023; Ram et al., 2023; Borgeaud et al., 2022; Khandelwal et al., 2019) have demonstrated success with zero or few-shot augmentation of standard LLM contexts with retrieved documents. Beyond context augmentation, several traditional NLP approaches (Lin et al., 2023b; Wang et al., 2023; Xu et al., 2023; Liu et al., 2024d; Xiang et al., 2024; Xu et al., 2024) have demonstrated that fine-tuning LLMs to be robust to noisy RAG outputs can lead to performance improvements.
Contemporary to VHs, Wang et al. (2024c) and Zhao et al. (2024) introduce NIAH-like synthetic benchmarks for multi-image and video QA applications, respectively. These datasets focus on short-range retrieval problems (<100 images/frames) and still consist of artificial content like out-of-distribution visual “needles.” In contrast, VHs uses realistic, in-distribution visual content in real images and scaling up to 10K images per query to pose a more complex, real-world challenge. Similarly, new benchmarks like CompBench (Kil et al., 2024), MANTIS (Jiang et al., 2024), and MUIRBench (Wang et al., 2024a) were recently proposed to evaluate LMMs’ general multi-image reasoning performance. However, these datasets are limited to fewer than 10 images per question and can make diagnosing specific LMM abilities difficult, as answering a single question often requires multiple intermixed capabilities. VHs distinguishes itself by offering a diagnostic, large-scale evaluation explicitly designed to test LMMs’ visual retrieval and cross-image reasoning capabilities without such confounding factors.